OpenStudy (anonymous):

How to find partial fraction decomposition??? ( x^2 - 7x + 27 ) / (x^2)(x^2 + 1)

OpenStudy (anonymous):

Just solve: \[\style{font-size:12px;font-family:'Courier New'}{\frac{x^2-7x+27}{x^2\left(x^2+1\right)}=\frac Ax+\frac B{x^2}+\frac{Cx+D}{x^2+1}}\] for A, B, C & D